OpenStudy (anonymous):

The leg of a right triangle is 2 units and the hypotenuse is 4 units. What is the length, in units, of the other leg of the triangle?

OpenStudy (anonymous):

Pythagorean Theorem: \(a^2 + b^2 = c^2\) Where 'a' and 'b' are the two legs, and 'c' is the hypotenuse. Plug in what we know: \(a^2 + 2^2 = 4^2\) Okay, what's 2 * 2?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

Yes, and what's 4 * 4?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

Yes, so we have: \(a^2 + 4 = 16\) Now we subtract 4 to both sides, what's 16 - 4?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

Yes, that gives us: \(a^2 = 12\) Now we have to find the square root of both sides: \(a = \sqrt{12}\)

OpenStudy (anonymous):

3.4641016

OpenStudy (anonymous):

Yes, so the length of the other leg is approximately 3.46.

OpenStudy (anonymous):

Thanks igreen for the help. I ask you for help because other people don't really help me with my problems! Thanks again!